# An Integrated Deep Learning Model for Skin Cancer Detection Using Hybrid Feature Fusion Technique

### v1

@@ 3.5 Feature Fusion Technique
Where $w_{\text{i}}$ is the weight value assigned to the score value of the $i$th model. According to the sensitivity analysis, the sum rule is the most durable method for error estimation. It has been found that the Sum rule performs better than most other state-of-the-art score-level fusion algorithms.
@@ 4.2 Evaluation Metrics
  • Accuracy Score: Accuracy measures how properly the classifier anticipates the classes 29. The average number of samples is correctly classified by the classifier. It is mathematically defined as:
@@ 5 Discussion
Table 3 presents a comparison of five state-of-the-art models proposed for skin cancer detection by different authors. M. S. Ali et al. 3 proposed a deep convolutional neural network (DCNN) model, achieving an accuracy of 91.93% on the HAM10000 dataset. K. Ali et al. 4 proposed an EfficientNetB4-based model, achieving an accuracy of 87.91% on the same dataset. D. Keerthan et al. 7 proposed an SVM-based model, achieving an accuracy of 88.2% on the ISBI2016 dataset. M. Abdar et al. 18 proposed a two-way deep belief network with deep learning (TWDBDL) model, achieving an accuracy of 88.95% on the Kaggle dataset.
@@
M. S. Ali et al. 3 2021 DCNN HAM10000 91.93%
K. Ali et al. 4 2022 EfficientNetB4 HAM10000 87.91%
D. Keerthan et al. 7 2023 SVM ISBI2016 88.2%
M. Abdar et al. 18 2021 TWDBDL Kaggle 88.95%
H.-Y. Huang et al. 30 2023 YOLOv5 ISIC 79.2%
Yuvika Gautam et al. 16 2024 FusionEXNet HAM10000 90.83%
Hatice Catal Reis et al. 31 2024 MABSCNET & ViT ISIC 78.63% & 76.50%
@@ 6 Conclusions
Acknowledgements.
The authors would like to extend their sincere appreciation to the Researchers Supporting Project Number (RSP2025R301), King Saud University, Riyadh, Saudi Arabia.

### v2

@@ 3.5 Feature Fusion Technique
Where $w_{\text{i}}$ is the weight value assigned to the score value of the $i$th model. According to the sensitivity analysis, the sum rule is the most durable method for error estimation. It has been found that the Sum rule performs better than most other state-of-the-art score-level fusion algorithms[29].
@@ 4.2 Evaluation Metrics
  • Accuracy Score: Accuracy measures how properly the classifier anticipates the classes [30]. The average number of samples is correctly classified by the classifier. It is mathematically defined as:
@@ 5 Discussion
Table 3 presents a comparison of five state-of-the-art models proposed for skin cancer detection by different authors. M. S. Ali et al. [3] proposed a deep convolutional neural network (DCNN) model, achieving an accuracy of 91.93% on the HAM10000 dataset. K. Ali et al. [4] proposed an EfficientNetB4-based model, achieving an accuracy of 87.91% on the same dataset. D. Keerthan et al. [7] proposed an SVM-based model, achieving an accuracy of 88.2% on the ISBI2016 dataset. M. Abdar et al. [18] proposed a two-way deep belief network with deep learning (TWDBDL) model, achieving an accuracy of 88.95% on the Kaggle dataset.
@@
M. S. Ali et al. [3] 2021 DCNN HAM10000 91.93%
K. Ali et al. [4] 2022 EfficientNetB4 HAM10000 87.91%
D. Keerthan et al. [7] 2023 SVM ISBI2016 88.2%
M. Abdar et al. [18] 2021 TWDBDL Kaggle 88.95%
H.-Y. Huang et al. [31] 2023 YOLOv5 ISIC 79.2%
Yuvika Gautam et al. [16] 2024 FusionEXNet HAM10000 90.83%
Hatice Catal Reis et al. [32] 2024 MABSCNET & ViT ISIC 78.63% & 76.50%
@@ 6 Conclusions
Declarations
  • Funding : This research was supported by the National Science and Technology for research fellowship of 2022-23.
  • Conflict of interest/Competing interests: The authors declare no conflict of interest.
  • Ethics approval and consent to participate: Not Applicable
  • Statement of human and animal rights: On behalf of all authors, the corresponding author affirms that human and animal rights were upheld in the study.
  • Consent statement: On behalf of all authors, the corresponding author states that informed consent was obtained from all participants involved in the study.
  • Data availability: The selected datasets are sourced from free and open-access source: https://www.kaggle.com/datasets/fanconic/skin-cancer-malignant-vs-benign?select=train
  • Materials availability: Not Applicable
  • Code availability: Not Applicable
  • Author contribution: Conceptualization: Maksuda Akter, Rabea Khatun, Md. Manowarul Islam; Supervision: Md. Manowarul Islam; Methodology: Maksuda Akter, Rabea Khatun, Md. Manowarul Islam,; Data curation: Maksuda Akter, Rabea Khatun; Resources: Maksuda Akter, Rabea Khatun, Md. Manowarul Islam; Software: Maksuda Akter, Rabea Khatun, Md. Manowarul Islam; Formal analysis, Maksuda Akter, Rabea Khatun, Md. Manowarul Islam, Md. Alamin Talukder; Writing – original draft: Maksuda Akter, Rabea Khatun; Validation, Md. Manowarul Islam, Dr. Md. Ashraf Uddin.